# Progressive-Resolution Policy Distillation: Leveraging Coarse-Resolution Simulations for Time-Efficient Fine-Resolution Policy Learning

### v2

@@ VII Discussions
Finer resolution assumption: In this paper, we assumed that finer resolution in excavation simulators leads to better transfer policy performance in real-world environments, supported by the experimental results in Section VI-F. However, this assumption may not always hold. The relationship between resolution and the reality gap could reverse or remain unchanged beyond a certain point. While finer resolutions may reveal physical differences, these might not significantly impact policy learning. The influence of policy generalization through DR on this relationship is also considered. These discussions are left for future work.

### v3

@@ VII Discussions
Finer resolution assumption: In this paper, we assumed that finer resolution in excavation simulators leads to better transfer policy performance in real-world environments, supported by the experimental results in Section VI-G. However, this assumption may not always hold. The relationship between resolution and the reality gap could reverse or remain unchanged beyond a certain point. While finer resolutions may reveal physical differences, these might not significantly impact policy learning. The influence of policy generalization through DR on this relationship is also considered. These discussions are left for future work.
@@
Simulation Enginnering Cost: In this paper, multi-resolution simulations are constructed as independent environments with different particle sizes, corresponding to varying spatial resolutions. Each resolution requires a separate simulation run with appropriately adjusted parameters to generate simulations at the corresponding resolution. We assume that such environments can be prepared in advance, and our framework is designed to efficiently learn high-resolution policies under this assumption. In practice, the engineering cost of creating variable-resolution simulators is relatively low in modern earthwork simulation platforms. For example, in particle-based simulators such as Vortex Studio [17] and OPERA [62], modifying a single configuration parameter is sufficient to control particle size and generate simulations at different resolutions.